OpenStudy (anonymous):

f(x) = x^2 - 4x - 3 is a function. How do I find x such that f(x) = 2?

OpenStudy (anonymous):

x^2 - 4x - 3 = 2 and solve

OpenStudy (anonymous):

I factor it?

OpenStudy (anonymous):

2=x^2-4x-3 that is x^2-4x-5=0 and factorize that

OpenStudy (anonymous):

which ever you prefear

OpenStudy (anonymous):

ohh, okay. thanks